OpenStudy (help!!!!):

The area of a square is 18 sq. in. less than twice the area of a triangle. The base of the triangle is 3 in longer than, and the altitude is the same length as, the side of the square. Find the length of the side of the square.

OpenStudy (help!!!!):

So far I got x^2=16 for the area of the square and 1/2x^2+3/2x for the area of the triangle. Please help! @jim_thompson5910 @mathstudent55 @mathmate @ganeshie8 @SolomonZelman

OpenStudy (mathmate):

 Let side of square = x then area of square = x^2 area of triangle = (1/2)x(x+3) the area of a square is 18 sq.in less than twice the area of a triangle means x^2+18=2(1/2)x(x+3) solve for x to get the answer for the side of the square.

OpenStudy (help!!!!):

@mathmate thank you so much for the answer and you time!

OpenStudy (mathmate):

You're welcome! :)